OpenStudy (kmullis6):

PLEASE HELP! Name the algebraic property demonstrated in the example below: 3 • (x • y) = (3 • x) • y Distributive Property Transitive Property Associative Property of Multiplication Commutative Property of Multiplication

OpenStudy (anonymous):

This is meant to indicate that order of multiplication is immaterial. It's the commutative property.

OpenStudy (kmeis002):

With all respect, these are easy to mix up. The commutative property is defined as the order of the operands dont matter: \[ x*y = y*x\] But the associative property is where the order of the operations dont matter: \[ x+(y+z) = (x+y)+z \]

OpenStudy (anonymous):

Thank you for catching that! @kmeis002